# Active Adaptive Experimental Design for Treatment Effect Estimation with Covariate Choices

### v1

@@ B.5 Derivation of the Semiparametric Efficiency Bound
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\frac{1}{T}\sum^{T}_{t=1}\left\{s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}\Bigg{]}$
$\ \ \ -\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y_{t}(0)\mid X]\big{)}}{w(0\mid X_{t})}\frac{q(X)}{p(X_{t})}\left\{s^{0}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\frac{1}{T}\sum^{T}_{t=1}\left\{s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}\Bigg{]}$
@@
$\ \ \ -\theta_{0}\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\Bigg{\{}\mathbbm{1}[A_{t}=1]\left\{s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right\}+\mathbbm{1}[A_{t}=0]\left\{s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right\}\Bigg{\}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{W}\Bigg{[}\left(\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right\}\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\right)\frac{1}{S}\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\Bigg{]},$
where we used $\mathbbm{1}[A_{t}=1]\mathbbm{1}[A_{t}=0]=0$. Here, it holds that
$\mathbb{E}_{W}\Bigg{[}\mathbbm{1}[A=d]\left\{s^{d}(Y(d)\mid X;\beta)+\frac{\dot{w}(d\mid X;\beta)}{w(d\mid X;\beta)}\right\}\mid X\Bigg{]}$
$=\mathbb{E}_{W}\Bigg{[}\mathbbm{1}[A=d]\mid X\Bigg{]}\left\{s^{d}(Y(d)\mid X;\beta)+\frac{\dot{w}(d\mid X;\beta)}{w(d\mid X;\beta)}\right\}$
@@
$\theta_{0}\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A=0]\big{(}Y-\mathbb{E}[Y_{t}(0)\mid X]\big{)}}{w(0\mid X_{t})}\right\}\frac{q(X)}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\left(\theta_{0}(\widetilde{X}_{s})-\theta_{0}\right)\Bigg{]}=0.$
@@
Here, from Lemma B.1, we have
@@
$=\mathbb{E}_{W}\Big{[}Y(1)s^{1}(Y(1)\mid X;\beta)-Y(0)s^{0}(Y(0)\mid X;\beta)+\theta(\widetilde{X};\beta)\widetilde{\zeta}(\widetilde{X};\beta)\Big{]},$
@@
Appendix C Proof of Theorem 4.1
@@
Appendix D Proof of Theorem 5.3
@@
=($=\left(\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)/\sqrt{\tau^{*}},$
@@
$\tau^{*}=\mathbb{E}_{X\sim p^{*}(x)}\left[\Bigg{\{}\frac{\sigma^{2}_{0}(1)(X)}{w^{*}(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w^{*}(0\mid X)}\Bigg{\}}\frac{q^{2}(X)}{p^{*2}(X)}\right].$
@@
=($=\left(\Big{\{}\mu(1)(X_{t})-\mu(0)(X_{t})-\widehat{\mu}_{t}(1)(X_{t})+\widehat{\mu}_{t}(0)(X_{t})\Big{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)/\sqrt{\tau^{*}}.$
@@
Proposition D.2 (Central Limit Theorem for an MDS; from Proposition 7.9, p. 194, Hamilton (1994); also see White (1984)).
@@
Under this proof strategy, we show Theorem 5.3 below. Our proof is inspired by Kato et al. (2020) and Kato et al. (2023). We denote $\mathbb{E}_{Y_{t}(a)\sim r^{a}(y\mid x),A_{t}\sim\widehat{w}_{t}(a\mid x),X_{t}\sim\widehat{p}_{t}(x)}$ by $\mathbb{E}_{P_{t}}$. In the proof, we use the following results.
Definition D.3.
@@
Proposition D.4 (Sufficient conditions for uniform integrability; Proposition 7.7, p. 191. Hamilton (1994)).
@@
Proposition D.5 ($L^{r}$ convergence theorem, p 165, Loeve (1977)).
@@
Step 1: check of condition (a).
@@
$\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}$
@@
Instead of $\sum^{T}_{t=1}\mathbb{E}_{P}[\xi 2_{t}(P)]$, we first consider the convergence of $\sum^{T}_{t=1}\mathbb{E}_{P}[(xi^{2}_{t}(P)|\mathcal{F}_{t-1}]=\Omega_{t}(P)$; that is, we show $\Omega_{t}(P)-1\xrightarrow{\mathrm{p}}0$. Then, by using the $L^{r}$-convergence theorem (Proposition D.5), we show $\sum^{T}_{t=1}\mathbb{E}_{P}[\xi^{2}_{t}(P)]-1\to 0$.
@@
$\mathbb{E}_{P_{t}}\left[\left(\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}\mid\mathcal{F}_{t-1}\right]$
$=\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}^{2}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}\mid\mathcal{F}_{t-1}\Bigg{]}$
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}$
$=\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}^{2}}{\widehat{w}^{2}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}^{2}}{\widehat{w}^{2}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}$
$=\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(X_{t})\big{)}^{2}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(X_{t})\big{)}^{2}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{\}}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\theta_{0}(X_{t})-\widehat{\theta}_{t}(X_{t})\Bigg{\}}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}.$

### v2

@@ A.5 Derivation of the Semiparametric Efficiency Bound
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\sum^{T}_{t=1}\mathbbm{1}[A_{t}=1]\left(s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right)\Bigg{]}$
$\ \ \ -\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X]\big{)}}{w(0\mid X_{t})}\frac{q(X)}{p(X_{t})}\left(s^{0}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\sum^{T}_{t=1}\mathbbm{1}[A_{t}=0]\left(s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)\Bigg{]}$
@@
$\ -\theta_{0}\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\left(\mathbbm{1}[A_{t}=1]\left(s^{1}(Y_{t}(1)\mid X_{t};\beta)+\frac{\dot{w}(1\mid X_{t};\beta)}{w(1\mid X_{t};\beta)}\right)+\mathbbm{1}[A_{t}=0]\left(s^{0}(Y_{t}(0)\mid X_{t};\beta)+\frac{\dot{w}(0\mid X_{t};\beta)}{w(0\mid X_{t};\beta)}\right)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{W}\Bigg{[}\left(\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right)\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\theta_{0}(\widetilde{X}_{s})\right)$
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\sum^{S}_{s=1}\widetilde{\zeta}(\widetilde{X}_{s};\beta)\Bigg{]},$
where we used $\mathbbm{1}[A_{t}=1]\mathbbm{1}[A_{t}=0]=0$.
Here, it holds that
$\mathbb{E}_{W}\Bigg{[}\mathbbm{1}[A=d]\left(s^{d}(Y(d)\mid X;\beta)+\frac{\dot{w}(d\mid X;\beta)}{w(d\mid X;\beta)}\right)\mid X\Bigg{]}$
$=\mathbb{E}_{W}\Bigg{[}\mathbbm{1}[A=d]\mid X\Bigg{]}\left(s^{d}(Y(d)\mid X;\beta)+\frac{\dot{w}(d\mid X;\beta)}{w(d\mid X;\beta)}\right)$
@@
$\mathbb{E}_{W}\Bigg{[}\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\mathbb{E}[Y_{t}(1)\mid X_{t}]\big{)}}{w(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\mathbb{E}[Y_{t}(0)\mid X_{t}]\big{)}}{w(0\mid X_{t})}\right)\frac{q(X_{t})}{p(X_{t})}+\frac{1}{S}\sum^{S}_{s=1}\left(\theta_{0}(\widetilde{X}_{s})-\theta_{0}\right)\Bigg{]}$
$=0.$
@@
Here, from Lemma A.1, we have
@@
$=\mathbb{E}_{W}\Big{[}Y(1)s^{1}(Y(1)\mid X;\beta)-Y(0)s^{0}(Y(0)\mid X;\beta)+\theta(\widetilde{X};\beta)\widetilde{\zeta}(\widetilde{X};\beta)\Big{]}.$
@@
Appendix B Proof of Theorem 4.1
@@
Appendix C Proof of Theorem 5.3
@@
$=\left(\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\right)\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)/\sqrt{\tau^{*}},$
@@
$\tau^{*}=\mathbb{E}_{X\sim p^{*}(x)}\left[\left(\frac{\sigma^{2}_{0}(1)(X)}{w^{*}(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w^{*}(0\mid X)}\right)\frac{q^{2}(X)}{p^{*2}(X)}\right].$
@@
$=\left(\left(\mu(1)(X_{t})-\mu(0)(X_{t})-\widehat{\mu}_{t}(1)(X_{t})+\widehat{\mu}_{t}(0)(X_{t})\right)\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)/\sqrt{\tau^{*}}.$
@@
Proposition C.2 (Central Limit Theorem for an MDS; from Proposition 7.9, p. 194, Hamilton (1994); also see White (1984)).
@@
Under this proof strategy, we show Theorem 5.3 below. Our proof is inspired by Kato et al. (2020). We denote $\mathbb{E}_{Y_{t}(a)\sim r^{a}(y\mid x),A_{t}\sim\widehat{w}_{t}(a\mid x),X_{t}\sim\widehat{p}_{t}(x)}$ by $\mathbb{E}_{P_{t}}$. In the proof, we use the following results.
Definition C.3.
@@
Proposition C.4 (Sufficient conditions for uniform integrability; Proposition 7.7, p. 191. Hamilton (1994)).
@@
Proposition C.5 ($L^{r}$ convergence theorem, p 165, Loeve (1977)).
@@
Step 1: check of condition (a)
@@
$\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\right)\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}$
@@
Instead of $\sum^{T}_{t=1}\mathbb{E}_{P}[\xi 2_{t}(P)]$, we first consider the convergence of $\sum^{T}_{t=1}\mathbb{E}_{P}[(xi^{2}_{t}(P)|\mathcal{F}_{t-1}]=\Omega_{t}(P)$; that is, we show $\Omega_{t}(P)-1\xrightarrow{\mathrm{p}}0$. Then, by using the $L^{r}$-convergence theorem (Proposition C.5), we show $\sum^{T}_{t=1}\mathbb{E}_{P}[\xi^{2}_{t}(P)]-1\to 0$.
@@
$\mathbb{E}_{P_{t}}\left[\left(\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\right)\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}+\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}\mid\mathcal{F}_{t-1}\right]$
@@
$=\mathbb{E}_{P_{t}}\Bigg{[}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\right)^{2}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\right)\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}$
$=\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}^{2}}{\widehat{w}^{2}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}^{2}}{\widehat{w}^{2}_{t}(0\mid X_{t})}\Bigg{)}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{)}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}\mid\mathcal{F}_{t-1}\Bigg{]}$
@@
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}$
$=\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(X_{t})\big{)}^{2}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(X_{t})\big{)}^{2}}{\widehat{w}_{t}(0\mid X_{t})}\Bigg{)}\frac{q^{2}(X_{t})}{\widehat{p}^{2}_{t}(X_{t})}\mid\mathcal{F}_{t-1}\Bigg{]}$
$\ \ \ +2\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\theta_{0}(X_{t})-\widehat{\theta}_{t}(X_{t})\Bigg{)}\frac{q(X_{t})}{\widehat{p}_{t}(X_{t})}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}\mid\mathcal{F}_{t-1}\Bigg{]}$
$\ \ \ +\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{)}^{2}\mid\mathcal{F}_{t-1}\Bigg{]}.$